# Characteristic Learning for Provable One Step Generation

### v3

@@ 1.4. Organization
The remainder of this article is organized as follows. Section 2 introduces the characteristic-driven generative learning method. A thorough analysis for this method is provided in Section 3. Section 4 presents numerical studies and discussions. Section 5 discusses related work and provides additional insights. Finally, Section 6 presents the conclusion and discusses future work. The proof of theoretical results can be found in the Appendix.
@@ 2. Characteristic Generative Learning
Dating back to Moser (1965); Dacorogna and Moser (1990), researchers proposed a continuous dynamic-induced approach for solving the normalizing equation (1.1). In the field of deep generative learning, flow-based models utilize ODE-dynamics to construct probability flows, effectively pushing-forward the prior distribution towards the target distribution. This family of generative models is represented by continuous normalizing flows (CNF) (Chen et al., 2018; Grathwohl et al., 2019) and their variants (Gao et al., 2019; Rozen et al., 2021; Gao et al., 2022; Lipman et al., 2023; Neklyudov et al., 2023; Albergo and Vanden-Eijnden, 2023). The major challenges faced by flow-based models revolve around two key questions:
@@ 2.1. Characteristics and Probability Flow ODE
    $\alpha(t)$ and $-\beta(t)$ are monotonically decreasing, and
@@
    $\dot{\alpha}(t),\ddot{\alpha}(t)\in C([0,1))$, $\dot{\alpha}(t)\alpha(t)\in C([0,1])$ and $\dot{\beta}(t),\ddot{\beta}(t)\in C([0,1])$.
In this paper, we focus on two examples shown in Table 2: linear interpolants and Föllmer flow. Both of them are widely-used in generative learning, such as Nichol and Dhariwal (2021); Liu et al. (2022); Albergo and Vanden-Eijnden (2023); Albergo et al. (2023a); Lipman et al. (2023); Chang et al. (2024).
@@
Linear interpolants (Liu et al., 2022; Lipman et al., 2023) $1-t$ $t$
@@
Denote by $\mu_{t}$ the distribution of the process $X_{t}$ for each $t\in(0,1)$. The following proposition demonstrates that $\mu_{t}$ has a density $\rho_{t}$ that interpolates between $\rho_{0}$ and $\rho_{1}$. Further, the density $\rho_{t}$ satisfies the continuous equation.
@@
The distribution of the stochastic interpolant $X_{t}$ has a density function $\rho(t,x)$ satisfies $\rho(0,x)=\rho_{0}(x)$, $\rho(1,x)=\rho_{1}(x)$, and
@@
As our primary objective is to generate samples that obey to the target distribution, we now consider the transport equation (2.2) from the lens of particles. It suffices to consider characteristics (Courant and Hilbert., 1989, Section II.2), along which the transport equation becomes an ODE
@@
where $x(t)$ is a characteristic, representing the position of particle at time $t\in(0,1)$, and $b^{*}$ is the associated velocity field that moves particles around. This characteristic ODE (Evans, 2010, Section 3.2) is known as the probability flow ODE (Song et al., 2021c). Further, the associated two-parameter continuous normalizing flow
$g_{t,s}^{*}:{\mathbb{R}}^{d}\rightarrow{\mathbb{R}}^{d},\quad x_{t}\mapsto x_{s},\quad 0\leq t\leq s\leq 1,$
where $x_{s}=x(s)$ represents the solution of (2.4) at time $s$ given $x(t)=x_{t}$. Notice that the flow $g_{t,s}^{*}$ pushes-forward the distribution $\mu_{t}$ onto $\mu_{s}$, that is,
$(g_{t,s}^{*})_{\sharp}\mu_{t}=\mu_{s},\quad 0\leq t\leq s<1.$
It is evident that the flow $g_{t,s}^{*}$ satisfies the semi-group property as follows.
Proposition 2.2 (Semi-group property).
@@
Observe that the flow $g^{*}$ satisfies the normalizing equation (1.1), and for each fixed $x_{t}$, $\{g(t,s,x_{t})\}_{s\geq t}$ is a part of the characteristic. Consequently, the generative learning can be reduced to the problem of estimate the characteristic $g^{*}(t,s,x)$, which minimizes the following quadratic risk
@@
where $Z_{t}=g_{0,t}^{*}(Z_{0})$, $Z_{s}=g_{0,s}^{*}(Z_{0})$ and $T\in(1/2,1)$ is a pre-specified stopping time.
@@ 2.2. Velocity Matching
(2.7) ^$\widehat{{\mathcal{L}}}_{n}(b)=\frac{1}{n}\sum_{i=1}^{n}\|\dot{\alpha}(t^{(i)})X_{0}^{(i)}+\dot{\beta}(t^{(i)})X_{1}^{(i)}-b(t^{(i)},X_{t}^{(i)})\|_{2}^{2},$
@@
7:     Compute the empirical risk ^$\widehat{{\mathcal{L}}}_{n}(b_{\theta})$ in (2.7).
@@
Proposition 2.3.
The 2-Wasserstein distance between the probability flow (2.8) and the estimated flow (2.9) at the stopping time $T$ is bounded as follows:
$W_{2}(\widehat{\mu}_{T},\mu_{T})\leq\exp(\|\widehat{b}\|_{\operatorname{Lip}}T)\int_{0}^{T}{\mathbb{E}}_{X_{t}\sim\mu_{t}}\Big{[}\|\widehat{b}(t,X_{t})-b^{*}(t,X_{t})\|_{2}^{2}\Big{]}\,\mathrm{d}t,$
where the Lipschitz constant of the velocity estimator $\widehat{b}$ is defined as
$\|\widehat{b}\|_{\operatorname{Lip}}=\sup\Big{\{}\|\nabla\widehat{b}(t,x)\|_{\operatorname{op}}:(t,x)\in(0,1)\times{\mathbb{R}}^{d}\Big{\}}.$
The error bound in Proposition 2.3 depends exponentially on the Lipschitz constant of the estimated velocity, which requires us to constrain the Lipschitz constant of the deep neural network. In practical applications, various techniques for restricting the Lipschitz constant of deep neural networks have been proposed, such as weight clipping (Arjovsky et al., 2017), gradient penalty (Gulrajani et al., 2017), spectral normalization (Miyato et al., 2018), and Lipschitz network (Zhang et al., 2022).
@@ 2.3. Euler Sampling
It is true that the Euler flow $\widehat{E}_{k,\ell}^{\tau}$ pushes-forward the distribution $\mu_{k\tau}$ approximately onto the distribution $\mu_{\ell\tau}$ for each $1\leq k\leq\ell\leq K$. Hence the Euler flow $\widehat{E}_{k,\ell}^{\tau}$ is an alternative approach for evaluating the flow $g_{t,s}^{*}$. However, it is important to note that the Euler sampling incurs a substantial computational cost, as it necessitates a large number of velocity network evaluations. This makes the Euler sampling encounter challenges in time-sensitive application scenarios. Therefore, there is a pressing need to develop an efficient simulation-free approach for evaluating the flow $g_{t,s}^{*}$.
@@ 2.4. Characteristic Fitting and Semi-Group Penalty
[…]
@@
Despite that both the continuous flow $g_{t,s}^{*}$ and Euler flow $\widehat{E}_{k,\ell}^{\tau}$ satisfy semi-group property (Propositions 2.2 and 2.4), the characteristic generator $\widehat{g}$ defined as (2.12) does not satisfy the semi-group property in general. In order to ensure the long-term stability of the estimator, we introduce the semi-group constraint, which requires
@@
(2.13) ^$\widehat{{\mathcal{R}}}_{m,K}^{\operatorname{Euler},\lambda}(g)=\widehat{{\mathcal{R}}}_{m,K}^{\operatorname{Euler}}(g)+\lambda\widehat{{\mathscr{P}}}(g),$
where $\lambda>0$ is the penalty parameter, and the semi-group-penalty is defined as
@@
The complete training and sampling procedures of the characteristic generator is summarized in Algorithms 23 and 4.
@@
12:     Compute the empirical risk ^$\widehat{{\mathcal{R}}}_{m,K,n}(g_{\phi})$ in (2.11) or (2.13).
@@ 3.1. Discussions of Assumptions
Assumption 1 is standard and commonly-used in flow-based generative models (Liu et al., 2022; Lipman et al., 2023; Albergo and Vanden-Eijnden, 2023). The following proposition demonstrates that the velocity field $b^{*}(t,x)$ is a spatial linear combination of $x$ and the score function $s^{*}(t,x)=\nabla\log\rho_{t}(x)$. This connection suggests that stochastic interpolants that satisfy Assumption 1 are closely related to diffusion models (Ho et al., 2020; Song et al., 2021c).
@@
Assumption 2 necessitates that the target distribution be a mixture of a Gaussian distribution and a distribution of compact support. Previous research (Lee et al., 2023; Oko et al., 2023; Chen et al., 2023a, d) has investigated the assumption of compact support in the target distribution. However, the probability flow that pushes a Gaussian prior distribution towards a target distribution with compact support may lack regularity because the velocity field is not generally Lipschitz. Therefore, it becomes crucial to impose additional intractable regularity properties on the probability flow ODE. In contrast, the probability flow towards a distribution of compact support with Gaussian smoothing exhibits high regularity even without any additional assumptions. The regularity properties of the probability flow ODE are established in Section 3.2. Furthermore, Assumption 2 covers a wide range of target distributions. In essence, our generative model has the capability to learn arbitrarily complex distributions with compact supports after Gaussian smoothing. This guarantees the application of our model to a variety of generative problems.
@@ 3.2. Properties of Probability Flow ODE
$\max_{1\leq k\leq d}|b_{k}^{*}(t,x)|\leq B_{\operatorname{vel}}R,\quad(t,x)\in(0,1)\times{\mathbb{B}}_{R}^{\infty},$
@@
$\max_{1\leq k\leq d}|g_{k}^{*}(t,s,x)|\leq B_{\operatorname{flow}}R,\quad 0\leq t\leq s\leq T,\leavevmode\nobreak\ x\in{\mathbb{B}}_{R}^{\infty},$
@@
$\max\Big{\{}\|\partial_{t}g^{*}(t,s,x)\|_{2},\|\partial_{s}g^{*}(t,s,x)\|_{2}\Big{\}}\leq B_{\operatorname{vel}}R,\quad 0\leq t\leq s\leq T,\leavevmode\nobreak\ x\in{\mathbb{B}}_{R}^{\infty}.$
@@
$\|\nabla b^{*}(t,x)\|_{\operatorname{op}}\leq G,\quad(t,x)\in(0,1)\times{\mathbb{R}}^{d},$
@@
$\|b^{*}(t,x)-b^{*}(t,x^{\prime})\|_{2}\leq G\|x-x^{\prime}\|_{2},\quad(t,x,x^{\prime})\in(0,1)\times{\mathbb{R}}^{d}\times{\mathbb{R}}^{d}.$
The uniform Lipschitz continuity of the velocity field or score plays a crucial role in controlling the error accumulation along the ODE or Euler method, as detailed in Corollary 3.10 and Theorem 3.11. Previous studies have often made the direct assumption of Lipschitz continuity for the velocity field or score at each time $t\in(0,1)$ (Chen et al., 2023d, a, c, b; Benton et al., 2024b; Gao and Zhu, 2024). However, this assumption appears to be restrictive, as it is difficult to verify. On the contrary, Assumptions 1 and 2 are easily satisfied and cover a large amount of generative tasks.
@@
$\|\nabla g^{*}(t,s,x)\|_{\operatorname{op}}\leq\exp(G(s-t)),\quad 0\leq t\leq s\leq T,\leavevmode\nobreak\ x\in{\mathbb{B}}_{R}^{\infty}.$
@@
$\|\partial_{t}b^{*}(t,x)\|_{2}\leq D\kappa(T)R:=D\sup_{t\in[0,T]}\Big{(}\frac{\dot{\alpha}_{t}^{2}}{\alpha_{t}^{2}}+\frac{|\ddot{\alpha}_{t}|}{\alpha_{t}}\Big{)}R,\quad(t,x)\in[0,T]\times{\mathbb{B}}_{R}^{\infty},$

### v5

@@ 1.4. Organization
The remainder of this article is organized as follows. Section 2 introduces the characteristic-driven generative learning method. A thorough analysis for this method is provided in Section 3. Section 4 presents numerical studies and discussions. Section 5 discusses related work and provides additional insights. Finally, Section 6 presents the conclusion and discusses future work. The theoretical proofs and experimental details are provided in the supplementary materials.
@@ 2. Characteristic Generative Learning
Dating back to Moser (1965); Dacorogna and Moser (1990), researchers proposed a continuous dynamic-induced approach for solving the normalizing equation (1.1). In the field of deep generative learning, flow-based models utilize ODE-dynamics to construct probability flows, effectively pushing the prior distribution towards the target distribution. This family of generative models is represented by continuous normalizing flows (CNF) (Chen et al., 2018; Grathwohl et al., 2019) and their variants (Gao et al., 2019; Rozen et al., 2021; Gao et al., 2022; Lipman et al., 2023; Neklyudov et al., 2023; Albergo and Vanden-Eijnden, 2023). The major challenges faced by flow-based models revolve around two key questions:
@@ 2.1. Characteristics and Probability Flow ODE
    $\alpha(t)$ and $\beta(t)$ are monotonically decreasing, and
@@
    $\dot{\alpha}(t),\ddot{\alpha}(t)\in C([0,1))$, $\dot{\alpha}(t)\alpha(t)\in C^{1}([0,1])$ and $\dot{\beta}(t),\ddot{\beta}(t)\in C([0,1])$.
In this paper, we focus on two examples shown in Table 3: linear interpolants and Föllmer flow. Both of them are widely used in generative learning, such as Nichol and Dhariwal (2021); Liu et al. (2022); Albergo and Vanden-Eijnden (2023); Albergo et al. (2023a); Lipman et al. (2023); Chang et al. (2024).
@@
Linear interpolants (Liu et al., 2022; Lipman et al., 2023) 1t1-t $t$
@@
Denote by $\mu_{t}$ the distribution of the process $X_{t}$ for each $t\in(0,1)$. The following proposition demonstrates that $\mu_{t}$ has a density $\rho_{t}$ that interpolates between $\rho_{0}$ and $\rho_{1}$. Further, the density $\rho_{t}$ satisfies the continuity equation.
@@
The distribution of the stochastic interpolant $X_{t}$ has a density function $\rho(t,x)$ satisfying $\rho(0,x)=\rho_{0}(x)$, $\rho(1,x)=\rho_{1}(x)$, and
@@
As our primary objective is to generate samples that obey the target distribution, we now consider the transport equation (2.2) from the lens of particles. It suffices to consider characteristics (Courant and Hilbert., 1989, Section II.2), along which the transport equation becomes an ODE:
@@
where $x(t)$ is knwon as the characteristic curve of the transport equation (2.2), representing the position of the particle at time $t\in(0,1)$, and $b^{*}$ is the associated velocity field that moves particles around. This characteristic ODE (Evans, 2010, Section 3.2) is known as the probability flow ODE (Song et al., 2021c).
Further, we define the associated two-parameter continuous normalizing flow $g_{t,s}^{*}$, which transports particles along the characteristic curves:
$g_{t,s}^{*}:{\mathbb{R}}^{d}\rightarrow{\mathbb{R}}^{d},\quad x_{t}\mapsto x_{s},\quad 0\leq t\leq s\leq 1.$
Here $x_{s}=x(s)$ represents the solution of (2.4) at time $s$ given $x(t)=x_{t}$. Notice that the flow $g_{t,s}^{*}$ pushes the distribution $\mu_{t}$ onto $\mu_{s}$, that is,
$(g_{t,s}^{*})_{\sharp}\mu_{t}=\mu_{s},\quad 0\leq t\leq s\leq 1.$
It is evident that the flow $g_{t,s}^{*}$ satisfies the semigroup property as follows.
Proposition 2.2 (Semigroup property).
@@
Observe that the flow map $g_{0,1}^{*}$ satisfies the normalizing equation (1.1), and for each fixed $x_{t}$, $\{g_{t,s}(x_{t})\}_{s\geq t}$ is a part of the characteristic curve. Consequently, the generative learning can be reduced to the problem of fitting the characteristic curves by minimizing the following quadratic risk:
@@
where $Z_{t}=g_{0,t}^{*}(Z_{0})$, $Z_{s}=g_{0,s}^{*}(Z_{0})$, and $T\in(1/2,1)$ is a pre-specified early-stopping time. The primary goal of our method, “characteristic learning”, is dedicated to estimating the flow map $g_{t,s}^{*}$ specified by the characteristic curves.
@@ 2.2. Velocity Matching
(2.7) $\widehat{{\mathcal{L}}}_{n}(b)=\frac{1}{n}\sum_{i=1}^{n}\|\dot{\alpha}(t^{(i)})X_{0}^{(i)}+\dot{\beta}(t^{(i)})X_{1}^{(i)}-b(t^{(i)},X_{t}^{(i)})\|_{2}^{2},$
@@
7:  Compute the empirical risk $\widehat{{\mathcal{L}}}_{n}(b_{\theta})$ in (2.7).
@@ 2.3. Euler Sampling
It is true that the Euler flow $\widehat{E}_{k,\ell}^{\tau}$ pushes the distribution $\mu_{k\tau}$ approximately onto the distribution $\mu_{\ell\tau}$ for each $1\leq k\leq\ell\leq K$. Hence, the Euler flow $\widehat{E}_{k,\ell}^{\tau}$ is an alternative approach for evaluating the flow $g_{t,s}^{*}$. However, it is important to note that the Euler sampling incurs a substantial computational cost, as it necessitates a large number of velocity network evaluations. This makes the Euler sampling encounter challenges in time-sensitive application scenarios. Therefore, there is a pressing need to develop an efficient simulation-free approach for evaluating the flow $g_{t,s}^{*}$.
@@ 2.4. Characteristic Fitting and Semi-Group Penalty
(2.10) $\widehat{{\mathcal{R}}}_{T,m,K}^{\operatorname{Euler}}(g)=\frac{2}{mK^{2}}\sum_{i=1}^{m}\sum_{k=0}^{K-1}\Bigg\{\frac{1}{2}\|\widehat{Z}_{k}^{(i)}-g(t_{k},t_{k},\widehat{Z}_{k}^{(i)})\|_{2}^{2}+\sum_{\ell=k+1}^{K-1}\|\widehat{Z}_{\ell}^{(i)}-g(t_{k},t_{\ell},\widehat{Z}_{k}^{(i)})\|_{2}^{2}\Bigg\}.$
@@
Despite that both the continuous flow $g_{t,s}^{*}$ and Euler flow $\widehat{E}_{k,\ell}^{\tau}$ satisfy semigroup property (Propositions 2.2 and 2.3), the characteristic generator $\widehat{g}$ defined as (2.11) does not satisfy the semi-group property in general. In order to ensure the long-term stability of the estimator, we introduce the semi-group constraint, which requires
@@
(2.12) $\widehat{{\mathcal{R}}}_{m,K}^{\operatorname{Euler},\lambda}(g)=\widehat{{\mathcal{R}}}_{m,K}^{\operatorname{Euler}}(g)+\lambda\widehat{{\mathscr{P}}}(g),$
where $\lambda>0$ is the penalty parameter, and the semigroup-penalty is defined as
@@
The complete training and sampling procedures of the characteristic generator are summarized in Algorithms 2, 3 and 4.
@@
12:  Compute the empirical risk $\widehat{{\mathcal{R}}}_{m,K,n}(g_{\phi})$ in (2.10) or (2.12).
@@ 3.1. Discussions of Assumptions
Assumption 1 is standard and commonly-used in flow-based generative models (Liu et al., 2022; Lipman et al., 2023; Albergo and Vanden-Eijnden, 2023). The following proposition demonstrates that the velocity field $b^{*}(t,x)$ is a spatial linear combination of $x$ and the score function $s^{*}(t,x)=\nabla\log\rho_{t}(x)$. This connection suggests that stochastic interpolants that satisfy Assumption 1 are closely related to score-based diffusion models (Ho et al., 2020; Song et al., 2021c).
@@
Assumption 2 necessitates that the target distribution be a Gaussian convolution of a distribution with compact support. Previous research (Lee et al., 2023; Oko et al., 2023; Chen et al., 2023a, d; Beyler and Bach, 2025) has investigated the assumption of compact support in the target distribution. However, the probability flow that pushes a Gaussian prior distribution towards a target distribution with compact support may lack regularity because the velocity field is not uniformly Lipschitz in general. Therefore, it becomes crucial to impose additional intractable regularity properties on the probability flow ODE. In contrast, the probability flow towards a distribution of compact support with Gaussian smoothing exhibits high regularity even without any additional assumptions. The regularity properties of the probability flow ODE are established in Section 3.2. Furthermore, Assumption 2 covers a wide range of target distributions, including Gaussian mixtures (Grenioux et al., 2024, Appendix C). In essence, our generative model is capable of learning arbitrarily complex distributions with compact support, provided they are smoothed via convolution with a Gaussian distribution of sufficiently small variance. This flexibility ensures that the model can be effectively applied to a wide range of generative tasks.
@@ 3.2. Properties of Probability Flow ODE
Suppose Assumptions 1 and 2 hold. Let $R\in(1,+\infty)$. Then it follows that
$\max_{1\leq k\leq d}|b_{k}^{*}(t,x)|\leq B_{\operatorname{vel}}R,\quad 0\leq t\leq 1,\penalty 10000\ x\in{\mathbb{B}}_{R}^{\infty},$
@@
Suppose Assumptions 1 and 2 hold. Let $R\in(1,+\infty)$. Then it follows that
$\max_{1\leq k\leq d}|g_{k}^{*}(t,s,x)|\leq B_{\operatorname{flow}}R,\quad 0\leq t\leq s\leq 1,\penalty 10000\ x\in{\mathbb{B}}_{R}^{\infty},$
@@
Suppose Assumptions 1 and 2 hold. Let $R\in(1,+\infty)$. Then it follows that
$\max\Big\{\|\partial_{t}g^{*}(t,s,x)\|_{2},\|\partial_{s}g^{*}(t,s,x)\|_{2}\Big\}\leq{B^{\prime}_{\operatorname{flow}}}R,\quad 0\leq t\leq s\leq 1,\penalty 10000\ x\in{\mathbb{B}}_{R}^{\infty}.$
where the constant $B^{\prime}_{\operatorname{flow}}$ only depends on $d$ and $\sigma$.
@@
$\|\nabla b^{*}(t,x)\|_{\operatorname{op}}\leq G,\quad 0\leq t\leq 1,\penalty 10000\ x\in{\mathbb{B}}_{R}^{\infty},$
@@
$\|b^{*}(t,x)-b^{*}(t,x^{\prime})\|_{2}\leq G\|x-x^{\prime}\|_{2},\quad 0\leq t\leq 1,\penalty 10000\ (x,x^{\prime})\in{\mathbb{R}}^{d}\times{\mathbb{R}}^{d}.$
The uniform Lipschitz continuity of the velocity field or score plays a crucial role in controlling the error accumulation along the ODE or Euler method, as detailed in Corollary 3.11 and Theorem 3.12. Previous studies have often made the direct assumption of Lipschitz continuity for the velocity field or score at each time $t\in(0,1)$ (Chen et al., 2023d, a, c, b; Benton et al., 2024b; Gao and Zhu, 2024; Fukumizu et al., 2025). However, this assumption appears to be restrictive, as it is difficult to verify. On the contrary, Assumptions 1 and 2 are easily satisfied and cover a large amount of generative tasks.
@@
$\|\nabla g^{*}(t,s,x)\|_{\operatorname{op}}\leq\exp(G(s-t)),\quad 0\leq t\leq s\leq 1,\penalty 10000\ x\in{\mathbb{R}}^{d}.$
@@
$\|\partial_{t}b^{*}(t,x)\|_{2}\leq D\kappa(T)R,\quad\kappa(T):=\sup_{t\in[0,T]}\Big(\frac{\dot{\alpha}_{t}^{2}}{\alpha_{t}^{2}}+\frac{|\ddot{\alpha}_{t}|}{\alpha_{t}}\Big),\quad(t,x)\in[0,T]\times{\mathbb{B}}_{R}^{\infty},$
@@
Remark 3.9.
Proposition 3.8 reveals a critical property of the velocity field: its time derivative, $\|\partial_{t}b(t,x)\|_{2}$, becomes unbounded as $t\rightarrow 1$. This behavior necessitates stopping the ODE integration at an endpoint $T<1$ due to two major consequences:
  1. (i)
    From the perspective of velocity approximation: A function with an exploding derivative is practically impossible for a neural network to approximate accurately near the singularity. This is a well-known challenge in flow-based and score-based modeling, as discussed in the literature (Kim et al., 2022; Pidstrigach, 2022; Duong et al., 2025).
  2. (ii)
    From the perspective of discretization: The error guarantees of numerical ODE solvers are contingent on the smoothness of the vector field. As $t\rightarrow 1$, the unbounded derivative violates these smoothness conditions, causing the solver’s discretization error to become uncontrolled.